OpenStudy (anonymous):

Combined Operations 2. Combine like terms. 7+4x-10x+15 A. 22-6x B. 11x-5x C. 22-14x D. 22+6x

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

Do you know what like terms are?

OpenStudy (anonymous):

can you help me i'll give you a medal and fan like

OpenStudy (anonymous):

they are terms that are alike

OpenStudy (anonymous):

Okay, can you point out two terms that are alike in the expression?

OpenStudy (anonymous):

um a

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

what you think

OpenStudy (anonymous):

Here is how I do it...

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[ \begin{split} 7+4x-10x+15 &= 7+15+4x-10x \\ &=22+(4-10)x \\ &=22-6x \end{split} \]

OpenStudy (anonymous):

i was right thank you

OpenStudy (anonymous):

So I factor the \(x\). Then you can add just fine.

OpenStudy (anonymous):

thanx can you help me again

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

and be my friend well fan

OpenStudy (anonymous):

more math problems

OpenStudy (anonymous):

Ok, create a new math question then

OpenStudy (anonymous):

ya